OpenStudy (anonymous):

from parade magazine, august 3,1997, page 8 by marilyn vass savant, can you help with this poblem? suppose it takes one man 5 hours to paint a house and it takes another man 3 hours to paint the same house. if the two men work together, how many hours would it take them? calculate the answer.

OpenStudy (anonymous):

heck no

OpenStudy (anonymous):

answer is \[\frac{5\times 3}{5+3}\]

OpenStudy (anonymous):

you need method or just the answer? it should be clear how i used the numbers and it is a very general solution if one man paint the house in \(x\) hours and the other in \(y\) hours then working together they will paint it in \(\frac{xy}{x+y}\) hours

OpenStudy (radar):

Unless they are painting overlapping areas lol

OpenStudy (radar):

by the way I like that product over sum, it works for parallel connected resistors when calculating total resistance.

OpenStudy (anonymous):

thank you satellite 73:)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@radar really?

OpenStudy (radar):

Yes, if there is more than two, you work them two at a time.

OpenStudy (anonymous):

oh! learn something new every day